alex is now 3 times as old as his sister mary. in 2 years alex will be 3 times mary's age then. how old was alex 4 years ago?
Answer by TimothyLamb(4379)

x = alex age
y = mary age
---
x = 3y
x + 2 = 3(y + 2)
---
put the system of linear equations into standard form
---
x - 3y = 0
x + 2 = 3y + 6
---
x - 3y = 0
x - 3y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
This linear system is singular. It has no solution
---
the reason the system has no solution is because there's an error in the problem statement:
x - 3y = 0
x - 3y = 4
it's impossible for (x - 3y) to equal 0 and 4
---
please check the problem statement and re-post your question.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---